When 702.5 J of heat energy is added to a 53.0 g sample of solid copper, the temperature increases to 87.6°C. What was the original temperature of the solid in °C? The specific heat capacity of solid copper is 0.385 J/g • K.
Answer #1

Heat = m×c×(∆T)

∆T = 702.5/53.0×0.385

(T-T​​​​i​​​) = 34.4 °c

T​​​​​​i = 87.6-34.4.

= 53.2 °c

So the original temperature was 53.2°c.